OpenStudy (anonymous):

Create an equation that results in at least one extraneous solution. Work through your equation, justify each step, and explain how the solution is extraneous.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

Ive never learned this. they never taught it..so can u explain please? @Hero

hero (hero):

Basically, you're given an equation, and asked to solve it, but when you solve it, you get two solutions, instead of one. Then upon checking the solutions, you realize one solution works and the other one doesn't. The other solution is called extraneous because it is not in the domain of variable x.

OpenStudy (anonymous):

So like x+3 = 10 +4?

OpenStudy (anonymous):

@Hero

hero (hero):

How is that extraneous? x + 3 = 10 + 4 x = 10 + 4 - 3 x = 11

hero (hero):

11 + 3 = 10 + 4 14 = 14

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

when you say the other solution doesnt work that means they dont equal?

OpenStudy (anonymous):

@Hero

hero (hero):

Yes, if you have an extraneous solution, then it will not work when you check it.